OpenStudy (anonymous):

The sum of one half of a number and 6 equals one third of the number, what is the number? a. -36 c. -24 b. -42 d. -48

OpenStudy (anonymous):

1/2x + 6 = 1/3x 6 = -1/6x 6 * -6 = -1/6x * -6 -36 = x check 1/2(-36) + 6 = 1/3(-36) -18 + 6 = -12 -12 = -12 so a. -36 :)

OpenStudy (anonymous):

thank you (:

OpenStudy (anonymous):

no problem! :)